# Stochastic Encodings for Active Feature Acquisition

### v1

@@ I.2 Model Specific Details
DIME.   DIME (Gadgil et al., 2024) uses two separate networks, one for prediction and one for predicting the CMI of features with the label. The information network is used to score each feature. The information network limits the output to a minimum of zero and maximum of the entropy of the current predictions $H(Y|\mathbf{x}_{S})$. The majority of the DIME implementation follows the GDFS implementation above. Instead of a temperature progression, we use an $\epsilon$ progression during main training following the original paper. This is given by $\epsilon$ Initial×$\times$[1.0, 0.25, 0.05, 0.005]. This gives the probability of choosing a feature uniformly at random compared to the best feature predicted by the information network. Main training is also done for 15 epochs for each $\epsilon$ value, the information network is trained to predict the change in loss when a given feature is acquired.
@@
  • Continuous Features. Continuous features first undergo a copula transformation $\tilde{x}_{i}=\Phi^{-1}(F_{i}(x_{i}))$, where $F_{i}$ is the empirical CDF of the feature and $\Phi$ is the standard normal CDF. This transformation is specific to stochastic encoders that regularize $I(X_{S};Z)$, since it enforces a symmetry associated with the mutual information and also encourages sparse disentangled representations (Wieczorek et al., 2018). We give $[m_{i}\tilde{x}_{i},m_{i}]$ to that feature’s encoder, which is given by the MLP described previously. The MLP ends with Batch Normalization which we found sped up training.
@@
  $b\sim\text{Bernoulli}(0.5)$
  $\mu_{f},\sigma_{f}=f_{\theta_{f}}^{e}(x_{f},m_{f}\times b)$
@@
Algorithm 2 Calculating the acquisition score for feature $i$.
@@
   $\text{Score}=\text{Score}+p(Y=y|\mathbf{x}_{O})\frac{1}{N_{2}}\frac{||\mathbf{g}_{\mathcal{G}_{i}}||_{2}}{\sum_{j}||\mathbf{g}_{\mathcal{G}_{j}}||_{2}}$
@@ 1 Acquisition Step 1 Training Step Syn 1 MiniBooNE MNIST Syn 1
ACFlow $\mathcal{O}(d)$ $\mathcal{O}(1)$ 1.483±0.0031.483$\pm$ 0.0031.483 ± 0.003 2.208±0.0022.208$\pm$ 0.0022.208 ± 0.002 21.318±0.19121.318$\pm$ 0.19121.318 ± 0.191 376.885±1.871376.885$\pm$ 1.871376.885 ± 1.871
DIME $\mathcal{O}(1)$ $\mathcal{O}(d)$ 0.017±0.0000.017$\pm$ 0.0000.017 ± 0.000 0.021±0.0000.021$\pm$ 0.0000.021 ± 0.000 0.024±0.0000.024$\pm$ 0.0000.024 ± 0.000 575.283±1.332575.283$\pm$ 1.332575.283 ± 1.332
EDDI $\mathcal{O}(d)$ $\mathcal{O}(1)$ 2.871±0.0592.871$\pm$ 0.0592.871 ± 0.059 9.039±0.2459.039$\pm$ 0.2459.039 ± 0.245 12.360±0.02912.360$\pm$ 0.02912.360 ± 0.029 273.358±1.060273.358$\pm$ 1.060273.358 ± 1.060
GDFS $\mathcal{O}(1)$ $\mathcal{O}(d)$ 0.018±0.0000.018$\pm$ 0.0000.018 ± 0.000 0.020±0.0000.020$\pm$ 0.0000.020 ± 0.000 0.024±0.0000.024$\pm$ 0.0000.024 ± 0.000 277.651±0.716277.651$\pm$ 0.716277.651 ± 0.716
GSMRL $\mathcal{O}(1)$ $\mathcal{O}(d)$ 0.047±0.0000.047$\pm$ 0.0000.047 ± 0.000 0.045±0.0000.045$\pm$ 0.0000.045 ± 0.000 0.097±0.0010.097$\pm$ 0.0010.097 ± 0.001 1010.650±3.1861010.650$\pm$ 3.1861010.650 ± 3.186
Opportunistic RL $\mathcal{O}(1)$ $\mathcal{O}(d)$ 0.029±0.0010.029$\pm$ 0.0010.029 ± 0.001 0.034±0.0010.034$\pm$ 0.0010.034 ± 0.001 0.030±0.0000.030$\pm$ 0.0000.030 ± 0.000 716.517±2.962716.517$\pm$ 2.962716.517 ± 2.962
VAE $\mathcal{O}(d)$ $\mathcal{O}(1)$ 0.153±0.0000.153$\pm$ 0.0000.153 ± 0.000 0.291±0.0010.291$\pm$ 0.0010.291 ± 0.001 0.452±0.0070.452$\pm$ 0.0070.452 ± 0.007 126.774±1.403126.774$\pm$ 1.403126.774 ± 1.403
SEFA (ours) $\mathcal{O}(|y|)$ $\mathcal{O}(1)$ 0.246±0.0010.246$\pm$ 0.0010.246 ± 0.001 0.314±0.0020.314$\pm$ 0.0020.314 ± 0.002 1.498±0.0021.498$\pm$ 0.0021.498 ± 0.002 297.061±1.701297.061$\pm$ 1.701297.061 ± 1.701

### v3

@@ I.2 Model Specific Details
[…]
@@
  • Continuous Features. Continuous features first undergo a copula transformation $\tilde{x}_{i}=\Phi^{-1}(F_{i}(x_{i}))$, where $F_{i}$ is the empirical CDF of the feature and $\Phi$ is the standard normal CDF. This transformation is specific to stochastic encoders that regularize $I(X_{S};Z)$, since it enforces a symmetry associated with the mutual information and also encourages sparse disentangled representations (Wieczorek et al., 2018). We give $m_{i}\tilde{x}_{i},m_{i}]$ to that feature’s encoder, which is given by the MLP described previously. The MLP ends with Batch Normalization which we found sped up training.
@@
  $u\sim\mathcal{U}(0,1)$
  $\mu_{f},\sigma_{f}=f_{\theta_{f}}^{e}(x_{f},m_{f}\times\mathcal{I}[u>p_{\text{Removal}}])$
@@
Algorithm 2 Calculating the acquisition score for feature ii.
@@
[…]
@@ 1 Acquisition Step 1 Training Step Syn 1 MiniBooNE MNIST Syn 1
ACFlow $\mathcal{O}(d)$ $\mathcal{O}(1)$ $1.483\pm 0.003$ $2.208\pm 0.002$ $21.318\pm 0.191$ $376.885\pm 1.871$
DIME $\mathcal{O}(1)$ $\mathcal{O}(d)$ $017\pm 0.000 0.0$$21\pm 0.000 0.0$$024\pm 0.000$ $575.283\pm 1.332$
EDDI $\mathcal{O}(d)$ $\mathcal{O}(1)$ $2.871\pm 0.059$ $9.039\pm 0.245$ $12.360\pm 0.029$ $273.358\pm 1.060$
GDFS $\mathcal{O}(1)$ $\mathcal{O}(d)$ $018\pm 0.000 0.0$$20\pm 0.000 0.0$$024\pm 0.000$ $277.651\pm 0.716$
GSMRL $\mathcal{O}(1)$ $\mathcal{O}(d)$ $047\pm 0.000 0.0$$45\pm 0.000 0.0$$097\pm 0.001$ $1010.650\pm 3.186$
Opportunistic RL $\mathcal{O}(1)$ $\mathcal{O}(d)$ $029\pm 0.001 0.0$$34\pm 0.001 0.0$$030\pm 0.000$ $716.517\pm 2.962$
VAE $\mathcal{O}(d)$ $\mathcal{O}(1)$ $0.153\pm 0.000$ $0.291\pm 0.001$ $0.452\pm 0.007$ $126.774\pm 1.403$
SEFA (ours) $\mathcal{O}(|y|)$ $\mathcal{O}(1)$ $0.246\pm 0.001$ $0.314\pm 0.002$ $1.498\pm 0.002$ $297.061\pm 1.701$